A 33-? resistor and a 25-? resistor are connected in parallel, and the combination is connected across a 280-V dc line.

(a) What is the resistance of the parallel combination?

---------- ?

(b) What is the total current through the parallel combination?

------------- A

(c) What is the current through each resistor?

I33? = -------------- A

I25? = -------------- A
Answer #1

a) Rnet = 33*25/(33+25)

= 14.22 ohms

b) I = V/Rnet

= 280/14.22

= 19.7 A

c) I33 = I*25/(33+25)

= 19.7*25/(33+25)

= 8.49 A

d) I25 = I*33/(25+33)

= 19.7*33/(33+25)

= 11.2 A
Answer #2

a)

1/R=1/33 + 1/25
R=14.22 ohms

b)

V=RI so I=V/R= 280/14.22=19.68 A

c)

I33=V/R: I=280/33=8.48 A
I25=V/R: I=280/25 = 11.2 A
Answer #3

Req = 33*25/33+25 = 14.22 ohm (ans)

Net current ,Ir=V/Req =280/14.22 =19.69 A (ans)

I33 =Ir*25/58 =8.49 A (ans)

I25 = Ir*33/58 = 11.20 A (ans)
Answer #4

1. Reff. = R1 * R2 / ( R1+R2)

= 25*33 / 58 = 14.22 OHM

2. v = I * Reff.

I = V/R = 280/14.22 = 19.69A

3. I1 = V/R1 = 280/33 = 8.48A

I2 = V/R2 = 280/25 = 11.2A